OpenStudy (anonymous):

Simplify 10 over 8-5i

OpenStudy (unklerhaukus):

\[\frac{10}{8-5i} = \frac{10}{8-5i}\times1\\ \qquad = \frac{10}{8-5i}\times\frac{8+5i}{8+5i}\\ \quad=\]